OpenStudy (anonymous):

Three hundred thirteen viewers were asked if they were satisfied with TV coverage of a recent disaster. Find P(satisfied | female). (Give your answer correct to two decimal places.) Find P(satisfied | male). (Give your answer correct to two decimal places.) Gender Female Male Satisfied 86 58 Not Satisfied 116 53

OpenStudy (anonymous):

\[P(\text{satisfied|male})=\frac{P(\text{satisfied}\cap\text{male})}{P(\text{male})}=\frac{\frac{58}{300}}{\frac{58+53}{300}}=\cdots\] You can figure out the other probability with similar reasoning.